# AdamX: An Adam improvement algorithm based on a novel exponential decay mechanism for the second-order moment estimate

### v1

@@ Abstract
Since the 21st century, artificial intelligence has been leading a new round of industrial revolution. Under the training framework, the optimization algorithm aims to stably converge high-dimensional optimization to local and even global minima. Entering the era of large language models, although the scale of model parameters and data has increased, Adam remains the mainstream optimization algorithm. However, compared with stochastic gradient descent (SGD) based optimization algorithms, Adam is more likely to converge to non-flat minima. To address this issue, the AdamX algorithm is proposed. Its core innovation lies in the proposition of a novel type of second-order moment estimation exponential decay rate, which gradually weakens the learning step correction strength as training progresses, and degrades to SGD in the stable training period, thereby improving the stability of training in the stable period and possibly enhancing generalization ability. Experimental results show that our second-order moment estimation exponential decay rate is better than the current second-order moment estimation exponential decay rate, and AdamX can stably outperform Adam and its variants in terms of performance. Our code is open-sourced at https://github.com/mengzhu0308/AdamX.
@@ 3.1 Impact of $\beta_{2}$ on the variance of the second-order moment estimation in Adam
$\bm{\xi}_{t}$ follows a normal distribution with zero mean and variance $\sigma^{2}$.
@@
Based on Hypothesis 1, Equation (1) can be written as:
@@
$\mathbb{E}[\bm{v}_{t}]=(1-\beta_{2}^{t})\sigma^{2}+(1-\beta_{2})\sum_{i=1}^{t}\beta_{2}^{t-i}\mathbb{E}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]$ (6)
@@
$\mathbb{E}[\overset{\frown}{\bm{v}}_{t}]=\mathbb{E}\left[\frac{\bm{v}_{t}}{1-\beta_{2}^{t}}\right]=\sigma^{2}+\frac{1-\beta_{2}}{1-\beta_{2}^{t}}\sum_{i=1}^{t}\beta_{2}^{t-i}\mathbb{E}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]$ (7)
@@
$=2\sigma^{4}(1-\beta_{2})^{2}\frac{1-\beta_{2}^{2t}}{1-\beta_{2}^{2}}+4\sigma^{2}(1-\beta_{2})^{2}\sum_{i=1}^{t}\beta_{2}^{2(t-i)}\mathbb{E}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]+$
$(1-\beta_{2})^{2}\sum_{i=1}^{t}\beta_{2}^{2(t-i)}\mathrm{Var}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]$
@@
$\lim\limits_{t\to\infty}\mathrm{Var}[\overset{\frown}{\bm{v}}_{t}]=\frac{1-\beta_{2}}{1+\beta_{2}}\left(2\sigma^{4}+4\sigma^{2}\mathbb{E}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]+\mathrm{Var}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]\right)$ (10)
Based on the analysis of Equation (10), during the training plateau phase (i.e., $t\to\infty$), the variance of the second moment estimate decreases as $\beta_{2}$ increases; if $\beta_{2}\to 1$, then $\lim\limits_{t\to\infty,\beta_{2}\to 1}\mathrm{Var}[\overset{\frown}{\bm{v}}_{t}]\to 0$, i.e., “zero fluctuation”, and Adam degenerates into momentum SGD. Whereas in Adam, $\beta_{2}$ is typically set to $0.999$, therefore Adam exhibits the “spike” phenomenon during the training plateau phase. To address this issue, the most straightforward approach is to set $\beta_{2}$ infinitely close to $1$. However, when $\beta_{2}$ is very close to $1$, the response of the second moment estimate to gradient changes becomes very slow. During training, the gradient distribution may change (e.g., from the early training stage to the plateau phase, or when the model encounters different data distributions), then $\overset{\frown}{\bm{v}}_{t}$ cannot adjust in time to adapt to the new gradient scale. Moreover, Adam uses the bias correction term $1/(1-\beta_{2}^{t})$ to counteract the zero initialization bias. When $\beta_{2}$ is very close to $1$, $\beta_{2}^{t}$ decays slowly, and the bias correction requires more iterations to take effect. In the early training stage, this causes $\overset{\frown}{\bm{v}}_{t}$ to underestimate the true second moment, resulting in excessively large update steps and increased instability in early training. Therefore, in Adam, $\beta_{2}$ cannot be too large nor too small. This is an irreconcilable contradiction.

### v2

@@ Abstract
Since the 21st century, artificial intelligence has been leading a new round of industrial revolution. Under the training framework, the optimization algorithm aims to stably converge high-dimensional optimization to local and even global minima. Entering the era of large language models, although the scale of model parameters and data has increased, Adam remains the mainstream optimization algorithm. However, compared with stochastic gradient descent (SGD) based optimization algorithms, Adam is more likely to converge to non-flat minima. To address this issue, the AdamX algorithm is proposed. Its core innovation lies in the proposition of a novel type of second-order moment estimation exponential decay rate, which gradually weakens the learning step correction strength as training progresses, and degrades to momentum SGD in the stable training period, thereby improving the stability of training in the stable period and possibly enhancing generalization ability. Experimental results show that our second-order moment estimation exponential decay rate is better than the current second-order moment estimation exponential decay rate, and AdamX can stably outperform Adam and its variants in terms of performance. Our code is open-sourced at https://github.com/mengzhu0308/AdamX.
@@ 3.1 Impact of $\beta_{2}$ on the variance of the second-order moment estimation in Adam
The full-batch gradient vector $\nabla f(\bm{\theta}_{t-1})$ and the noise vector $\bm{\xi}_{t}$ have independent and identically distributed (i.i.d.) components across dimensions.  [^1] This assumption was implicitly used in the derivations of arXiv v1 but was omitted from the formal statement of assumptions.
@@
Based on Hypothesis 2, Equation (1) can be written as:
@@
$\mathbb{E}[\bm{v}_{t}]=(1-\beta_{2}^{t})\sigma^{2}\bm{1}+(1-\beta_{2})\sum_{i=1}^{t}\beta_{2}^{t-i}\mathbb{E}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]$ (6)
@@
$\mathbb{E}[\overset{\frown}{\bm{v}}_{t}]=\mathbb{E}\left[\frac{\bm{v}_{t}}{1-\beta_{2}^{t}}\right]=\sigma^{2}\bm{1}+\frac{1-\beta_{2}}{1-\beta_{2}^{t}}\sum_{i=1}^{t}\beta_{2}^{t-i}\mathbb{E}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]$ (7)
@@
$=(1-\beta_{2})^{2}\sum_{i=1}^{t}\beta_{2}^{2(t-i)}\left(2\sigma^{4}\bm{1}+4\sigma^{2}\mathbb{E}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]+\mathrm{Var}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]\right)$
@@
$\lim\limits_{t\to\infty}\mathrm{Var}[\overset{\frown}{\bm{v}}_{t}]=\frac{1-\beta_{2}}{1+\beta_{2}}\left(2\sigma^{4}\bm{1}+4\sigma^{2}\mathbb{E}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]+\mathrm{Var}\left[(\nabla f(\bm{\theta}_{i-1}))^{2}\right]\right)$ (10)
Based on the analysis of Equation (10), during the training plateau phase (i.e., $t\to\infty$), the variance of the second moment estimate decreases as $\beta_{2}$ increases; if $\beta_{2}\to 1$, then $\lim\limits_{t\to\infty,\beta_{2}\to 1}\mathrm{Var}[\overset{\frown}{\bm{v}}_{t}]\to\bm{0}$, i.e., “zero fluctuation”, and Adam degenerates into momentum SGD. Whereas in Adam, $\beta_{2}$ is typically set to $0.999$, therefore Adam exhibits the “spike” phenomenon during the training plateau phase. To address this issue, the most straightforward approach is to set $\beta_{2}$ infinitely close to $1$. However, when $\beta_{2}$ is very close to $1$, the response of the second moment estimate to gradient changes becomes very slow. During training, the gradient distribution may change (e.g., from the early training stage to the plateau phase, or when the model encounters different data distributions), then $\overset{\frown}{\bm{v}}_{t}$ cannot adjust in time to adapt to the new gradient scale. Moreover, Adam uses the bias correction term $1/(1-\beta_{2}^{t})$ to counteract the zero initialization bias. When $\beta_{2}$ is very close to $1$, $\beta_{2}^{t}$ decays slowly, and the bias correction requires more iterations to take effect. In the early training stage, this causes $\overset{\frown}{\bm{v}}_{t}$ to underestimate the true second moment, resulting in excessively large update steps and increased instability in early training. Therefore, in Adam, $\beta_{2}$ cannot be too large nor too small. This is an irreconcilable contradiction.